OpenStudy (abbycross167):

Can someone please help me answer this question? Graph the following equations using the point-slope method. Graph the point from the equation in yellow, your other points in red and your line in blue. y - 7 = -3(x + 2)

OpenStudy (abbycross167):

@flcon04 Could you show me how to do this?

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (abbycross167):

well I did it once but my teacher said I got it wrong, would you like me to show you what I did?

OpenStudy (skullpatrol):

Ok

OpenStudy (abbycross167):



OpenStudy (abbycross167):

Sorry it took so long! but the dot on 1 is yellow and the dot on -3 is red

OpenStudy (abbycross167):

@skullpatrol

OpenStudy (skullpatrol):

the "point-slope method" First you need to change the given equation y-7=-3(x+2) into the point-slope form Recall the distributive property a(b+c)=ab+ac and use it on the right hand side of the equation. y-7=-3x-6 $$\huge y=-3x+1$$ Try to graph that^ :-)

OpenStudy (abbycross167):

isn't that what I did?

OpenStudy (skullpatrol):

Nope

OpenStudy (abbycross167):

oh

OpenStudy (skullpatrol):

You have the right y-intercept, but the slope is not right.

OpenStudy (abbycross167):

what would the slope be?

OpenStudy (skullpatrol):

-3

OpenStudy (abbycross167):

I'm confused, could you show me how to graph it? I really thought I did it right

OpenStudy (skullpatrol):

y = mx + b y = -3x + 1

OpenStudy (abbycross167):

I put a dot on -3

OpenStudy (skullpatrol):

Do you see why the slope m = -3?

OpenStudy (skullpatrol):

The dot you put on -3 is a point on the x-axis. That is not what is meant by the word "slope."

OpenStudy (abbycross167):

so -3 needs to be on te y-axis?

OpenStudy (skullpatrol):

Nope.

OpenStudy (skullpatrol):

Slope tells you how to go from one point on the line to another.

OpenStudy (abbycross167):

hmm I don't understand this

OpenStudy (skullpatrol):

Have you tried the khan academy on YouTube?

OpenStudy (abbycross167):

yes sir but I couldn't find a video that helped me with this problem... Could you please show me what the graph should look like?

OpenStudy (abbycross167):

Sir/ma'am*

OpenStudy (skullpatrol):

Start at the point (0,1) and go down 3 and to the right 1. That^ will give you a second point.

OpenStudy (skullpatrol):

(1, -2)

OpenStudy (abbycross167):



OpenStudy (abbycross167):

like that?

OpenStudy (skullpatrol):

Nope

OpenStudy (skullpatrol):

(0, 1), (1, -2)

OpenStudy (abbycross167):

oh I'm sorry, math isn't my best subject

OpenStudy (skullpatrol):

Plot those^ two points and draw a line through them :-)